OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 32x2 - 144

OpenStudy (anonymous):

2i, 12, -12 2i, 6i, -6i 2i, 6, -6 2i, 12i, -12i

OpenStudy (anonymous):

please help

OpenStudy (helder_edwin):

u can factor it like this \[\large f(x)=x^4-32x^2-144=(x^2-36)(x^2+4) \]

OpenStudy (helder_edwin):

so when u put f(x)=0 u have \[\large x^2-36=0\qquad\text{or}\qquad x^2+4=0\] i am sure u can finish this